OpenStudy (anonymous):

the perimeter of a rectangle with lenth 1and with w is given by p=2w+21 if the perimeter of a certain rectangle is 105 feet and the length is 35 feet what is the width

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

105=2w+2(35) => w=35/2=17.5

OpenStudy (anonymous):

how did you figure that out or could you show me steps to take when figuring this out thanks

OpenStudy (anonymous):

plug 35 for l (length), then solve for w: 105 = 2w + 70 2w = 35 w=

OpenStudy (anonymous):

see the equation is given by p=2w+2l, wer l is the length and w is the breadth....so for finding l given w we need to substitute it in out equation , as p=2*w+2(35) as length =35 is given in the question....so i guess u can figure out from here....right?

OpenStudy (anonymous):

sorry i had to step out for a minute yes i can figure it out from there thanks